The following are sales revenues for a large utility company for years 1 through 11. Forecast revenue for years 12 through 15. Because we are forecasting four years into the future, you will need to use linear regression as your forecasting method. (Enter your answers in millions.)

YEAR REVENUE (MILLIONS)
1 $ 4,873.2
2 5,070.1
3 5,522.8
4 5,720.8
5 5,507.2
6 5,193.6
7 $ 5,087.1
8 5,114.4
9 5,554.4
10 5,737.0
11 5,858.8

Period: Forecast:

12

13

14

15

16
Answer #1


PERIOD (X)

DEMAND (Y)

X

Y

X * Y

X^2

1

4873.2

1

4873.2

4873.2

1

2

5070.1

2

5070.1

10140.2

4

3

5522.8

3

5522.8

16568.4

9

4

5720.8

4

5720.8

22883.2

16

5

5507.2

5

5507.2

27536

25

6

5193.6

6

5193.6

31161.6

36

7

5087.1

7

5087.1

35609.7

49

8

5114.4

8

5114.4

40915.2

64

9

5554.4

9

5554.4

49989.6

81

10

5737

10

5737

57370

100

11

5858.8

11

5858.8

64446.8

121

SIGMA

66

59239.4

361493.9

506

INTERCEPT = (SIGMA(Y) * SIGMA(X^2) - SIGMA(X) * SIGMA(XY)) / (N * SIGMA(X^2) - SIGMA(X)^2)

INTERCEPT = (59239.4 * 506) - (66 * 361493.9) / ((11 * 506) - 66^2) = 5054.99

SLOPE = ((N * SIGMA(XY)) - (SIGMA(X) * SIGMA(Y))) - (N * SIGMA(X^2) - SIGMA(X)^2)

SLOPE = ((11 * 361493.9) - (66 * 59239.4) / ((11 * 506) - 66^2) = 55.07

LINE EQUATION = A + B(x), WHERE A IS THE INTERCEPT, B IS THE SLOPE, x IS THE PERIOD = 5054.99 + (55.07 * X)

FOR THE VALUE OF X = 12 FORECAST = 5054.99 + (55.07 * 12) = 5715.83

FOR THE VALUE OF X = 13 FORECAST = 5054.99 + (55.07 * 13) = 5770.9

FOR THE VALUE OF X = 14 FORECAST = 5054.99 + (55.07 * 14) = 5825.97

FOR THE VALUE OF X = 15 FORECAST = 5054.99 + (55.07 * 15) = 5881.04

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.